# Toward Patient-specific Partial Point Cloud to Surface Completion for Pre- to Intra-operative Registration in Image-guided Liver Interventions

### v1

@@ 2.2.1 Vector Neuron Occupancy Network (VN-OccNet):
To accomplish the rotation-equivariant point cloud completion task, VN-OccNet [1] leverages the original OccNet architecture [10], with a vector-based equivariant encoder. The decoder is invariant, where all the operations are non-vector. The target point cloud $\mathcal{T}=\{{t_{i}}\}_{i=1}^{N}\in\mathbb{R}^{3}$ is input to the encoder network, and outputs equivariant latent vector-list features ${z}$. The query set $\mathcal{X}=\{{x}_{i}\}_{i=1}^{K}\in\mathbb{R}^{3}$, set of points that encloses the liver, and latent feature ${z}$ are input to the decoder network to output the occupancy probabilities between 0 and 1. Points with probability greater than or equal to a threshold $c$ correspond to occupancy points that implicitly represent the liver surface. Since VN-OccNet aims to predict whether or not each point in the query point set corresponds to a liver surface point, binary cross-entropy classification loss is utilized to learn the parameters $\theta$.
@@ 3.2.2 Rotation Setup:
VN-OccNet is trained in three different modes based on the rotation imposed on the target point clouds: without any imposed rotations, with rotations restricted to the Z-axis, and with SO(3) rotations. We imposed random rotations within the range $[-\frac{\pi}{2},\frac{\pi}{2}]$ for Z-axis and SO(3) rotation modes. In the non-rotation mode, all target point clouds are aligned and share the same pose. The evaluation involves four train/test rotation settings: I/I, where both training and testing are conducted on non-rotated target point clouds; Z/Z, where both training and testing involve rotations of target point clouds along the Z-axis; Z/SO(3), where training is performed with Z-axis rotations, while testing involves SO(3) rotations; and SO(3)/SO(3), where both training and testing involve SO(3) rotations.
@@ 3.4 Registration Evaluation Setup
In a manner similar to the five target point clouds generated for each liver model during testing in the previous subsection 3.2, five target point clouds are generated for each in vitro phantom (No. 1 and No. 3) to perform registration. A set of 100 random SO(3) rotations within the range $[-\frac{\pi}{2},\frac{\pi}{2}]$ are applied to the target point clouds, resulting in 500 target point clouds at different orientations for each phantom. Registration w/ surface completion is performed by utilizing the complete target point clouds reconstructed from target point clouds featuring a minimum of 30% visibility. To test the influence of the increment of surface coverage, registration w/o surface completion was performed at three different visibilities of the target point cloud at around $30\%$, $40\%$, and $50\%$ by varying the value of $D$, where $D=1000,750,\text{and }600$, respectively.
@@ 3.5 Evaluation Metrics
where $|\mathcal{\hat{T}}|$ and $|\mathcal{G}|$ represent the total number of points in the complete generated target and ground truth target point cloud, respectively, and $d_{CD}(\mathcal{\hat{T}},\mathcal{G})$ denotes the Chamfer Distance between the point sets $\mathcal{\hat{T}}$ and $\mathcal{G}$.
@@
where $\text{occ}_{\text{gt}}$ and $\text{occ}_{\text{pred}}$ represent the ground truth and predicted occupancy values, respectively, and ||$\cdot$|| ⋅ | denotes the number of true occupancy labels.
@@ 4.2 Go-ICP Registration Results
Phantom W/O surface completion W/ surface completion
@@
Figure 6: Visualization of target fiducial markers (purple) and registered source fiducial markers (blue) for in vitro phantom No 1 and No 3. The target point clouds utilized for registration w/o surface completion have $30\%$ visibility, and the full point clouds are reconstructed from the same targets for registration w/ surface completion. The ground truth target liver surface is overlaid for improved visualization.

### v2

@@ 2.2.1 Vector Neuron Occupancy Network (VN-OccNet):
To accomplish the rotation-equivariant point cloud completion task, VN-OccNet [1] leverages the original OccNet architecture [10], with a vector-based equivariant encoder. The decoder is invariant, where all the operations are non-vector. The target point cloud $\mathcal{T}=\{{t_{i}}\}_{i=1}^{N}\in\mathbb{R}^{3}$ is input to the encoder network, and outputs equivariant latent vector-list features z{z}. The query set $\mathcal{X}=\{{x}_{i}\}_{i=1}^{K}\in\mathbb{R}^{3}$, set of points that encloses the liver, and latent feature z{z} are input to the decoder network to output the occupancy probabilities between 0 and 1. Points with probability greater than or equal to a threshold $c$ correspond to occupancy points that implicitly represent the liver surface. Since VN-OccNet aims to predict whether or not each point in the query point set corresponds to a liver surface point, binary cross-entropy classification loss is utilized to learn the parameters $\theta$.
@@ 3.2.2 Rotation Setup:
VN-OccNet is trained in three different modes based on the rotation imposed on the target point clouds: without any imposed rotations, with rotations restricted to the Z-axis, and with SO(3) rotations. We imposed random rotations within the range $\frac{\pi}{2},\frac{\pi}{2}]$ for Z-axis and SO(3) rotation modes. In the non-rotation mode, all target point clouds are aligned and share the same pose. The evaluation involves four train/test rotation settings: I/I, where both training and testing are conducted on non-rotated target point clouds; Z/Z, where both training and testing involve rotations of target point clouds along the Z-axis; Z/SO(3), where training is performed with Z-axis rotations, while testing involves SO(3) rotations; and SO(3)/SO(3), where both training and testing involve SO(3) rotations.
@@ 3.4 Registration Evaluation Setup
In a manner similar to the five target point clouds generated for each liver model during testing in the previous subsection 3.2, five target point clouds are generated for each in vitro phantom (No. 1 and No. 3) to perform registration. A set of 100 random SO(3) rotations within the range $\frac{\pi}{2},\frac{\pi}{2}]$ are applied to the target point clouds, resulting in 500 target point clouds at different orientations for each phantom. Registration w/ surface completion is performed by utilizing the complete target point clouds reconstructed from target point clouds featuring a minimum of 30% visibility. To test the influence of the increment of surface coverage, registration w/o surface completion was performed at three different visibilities of the target point cloud at around $30\%$, $40\%$, and $50\%$ by varying the value of $D$, where $D=1000,750,\text{and }600$, respectively.
@@ 3.5 Evaluation Metrics
[…]
@@
where $\text{occ}_{\text{gt}}$ and $\text{occ}_{\text{pred}}$ represent the ground truth and predicted occupancy values, respectively, and $|\cdot|$ denotes the number of true occupancy labels.
@@ 4.2 Go-ICP Registration Results
Phantom W/O surface completion
W/ surface
completion
@@
Figure 6: Visualization of target fiducial markers (blue) and registered source fiducial markers (purple) for in vitro phantom No 1 and No 3. The target point clouds utilized for registration w/o surface completion have $30\%$ visibility, and the full point clouds are reconstructed from the same targets for registration w/ surface completion. The ground truth target liver surface is overlaid for improved visualization.